OpenStudy (anonymous):

need help graphing x+4y=20

OpenStudy (anonymous):

get y by itself. y= 20-x/4, then graph

OpenStudy (anonymous):

(-4, 6), (-3, 5.75), (-2, 5.5), (0,5), (2, 4.5)

OpenStudy (anonymous):

Those are the coordinates that should draw that graph.

OpenStudy (dumbcow):

get y by itself to get equation is slope-intercept form 4y = -x +20 y = -x/4 +5 slope = -1/4 y-intercept = 5 A short cut to get 2 points fast is to plot intercepts plug in x=0, y=5 plug in y=0, x = 20 then draw line connecting 2 intercepts